# Unraveling Indirect In-Context Learning Using Influence Functions

### v1

@@ Abstract
This work introduces a novel paradigm for generalized In-Context Learning (ICL), termed Indirect In-Context Learning. In Indirect ICL, we explore demonstration selection strategies tailored for two distinct real-world scenarios: Mixture of Tasks and Noisy Demonstrations. We systematically evaluate the effectiveness of Influence Functions (IFs) as a selection tool for these settings, highlighting the potential for IFs to better capture the informativeness of examples within the demonstration pool. For the Mixture of Tasks setting, demonstrations are drawn from 28 diverse tasks, including MMLU, BigBench, StrategyQA, and CommonsenseQA. We demonstrate that combining BertScore-Recall (BSR) with an IF surrogate model can significantly improve performance, leading to average absolute accuracy gains of 0.37% and 1.45% for 3-shot and 5-shot setups when compared to traditional ICL metrics. In the Noisy Demonstrations setting, we examine scenarios where demonstrations might be mislabeled. Our experiments show that reweighting traditional ICL selectors (BSR and Cosine Similarity) with IF-based selectors boosts accuracy by an average of 2.90% for Cosine Similarity and 2.94% for BSR on noisy GLUE benchmarks. In sum, we propose a robust framework for demonstration selection that generalizes beyond traditional ICL, offering valuable insights into the role of IFs for Indirect ICL.
@@ 1 Introduction
Figure 1: An example showcasing demonstration selection for Indirect ICL using Influence Functions (IFs). Consider web corpora with many tasks (different from the end-task) and noisy data– Indirect ICL can be formalized as: Mixture of Tasks (Section 3.1) and Noisy (Section 3.2) ICL, respectively. In MoT, for a given target task (e.g. Medical Genetics), we first filter from this (indirect) pool of candidate demonstrations using BertScore and Cosine Similarity, then re-rank with IFs to select suitable demonstrations (e.g. High-School Biology).
Traditional ICL makes numerous assumptions that restrict its applicability to real-world problem domains. For instance, traditional ICL Min et al. (2021); Conneau (2019); Halder et al. (2020) assumes that demonstrations to be selected are directly and accurately annotated for the end-task. However, this is not always the case – for low-resource, sparse, or specialized domains, end-task information and labeled demonstrations might not be available. Similarly, when LLMs are deployed as services, the user query or the end task itself could be unknown beforehand, let alone providing direct demonstrations at inference. Thus, in this paper, we explore a more generalized setting for ICL, which we refer to as Indirect ICL. In indirect ICL, we aim to provide indirect (or incidental) supervision Yin et al. (2023); Li et al. (2024) by selecting demonstrations from a pool of examples where the majority are not directly suited to the end task due to severe distribution shifts. This includes selecting demonstrations from a pool that predominantly consists of examples belonging to other tasks, with few demonstrations from the end task possibly included. Additionally, the demonstration set may be mislabeled by humans Yan et al. (2014); Zhu et al. (2022) or LLMs Wu et al. (2023). Since the effectiveness of ICL heavily relies on the quality of demonstrations selected Kossen et al. (2024); Wu et al. (2022); Wang et al. (2024), selecting the most helpful indirect demonstrations becomes imperative in these situations.
@@
Traditional (direct) ICL methods that use metrics such as BertScore-Recall (BSR; Gupta et al. 2023a) and cosine similarity Reimers (2019) inherently rely on the semantic similarity between demonstrations and test samples. In this paper, we posit that IFs can be a reasonable measure of affinity between the end task and any (indirect) demonstrations. We show that it is practical to use IF to identify candidate demonstrations that represent a close inductive bias with the end that, and utilize this information for highly accurate demonstration selection in the challenging indirect ICL setting. As our experiments and results will demonstrate, this is indeed the case, and we find that IFs can aid in improved performance when simple semantic similarity is insufficient for demonstration selection.
@@
  • This work formalizes a new and general paradigm for ICL, namely Indirect In-Context Learning, where we benchmark demonstration selection for two distinct and real-world settings: (a) Mixture of Tasks and (b) Noisy Demonstrations.
  • We propose the use of Influence Functions (IFs) as a viable approach for demonstration selection in these generalized ICL settings.
  • For Mixture of Tasks, combining an IF Surrogate model with BertScore-Recall (BSR) can lead to a 0.37 and 1.45 average absolute increase in performance for $k=3$ and $k=5$ shots compared to the best performing traditional ICL metric.
  • For Noisy Demonstrations, we observe that undertaking a weighted average selection using traditional ICL selectors (BSR and Cosine Similarity) and IF based selectors increases the absolute accuracy on average by 2.90% for Cosine and 2.94% for BSR.
@@ 2.1 Traditional In-Context Learning
2.2 Indirect In-Conext Learning
@@
Noisy Demonstrations. To further generalize the problem of indirect ICL, we also consider noisy supervision that is likely existing in the pool of demonstrations. Formally, let $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ represent the training dataset, where $x_{i}\in X$ is the input and $y_{i}\in Y$ is the corresponding binary label. We randomly select 20% of the data points from $D$ and flip their labels.
Once the noisy dataset is generated, we use it for ICL. Formally, given the noisy set of input-output pairs $\{(x_{i},y_{i})\}_{i=1}^{k}$ and a test input $x_{\text{test}}$, the goal is to conditionally generate the test output $y_{\text{test}}$ based on the noisy training data.
@@ 2.3 Influence Functions
Recent work has also shown that theoretically, ICL is akin to implicit fine-tuning and that the transformer attention mechanism is actually undertaking (meta) gradient descent Dai et al. (2023). This implies that influence computation can serve as a beneficial tool to assess the usefulness of the demonstration set samples similar to the fine-tuning set.
@@ 3.1 Selecting within Mixture of Tasks
  • A pretrained-gradient based method where the samples are passed through the LLM itself. We then compute the IF using the extracted gradients.
Formally, for each validation example $(x_{\text{val}},y_{\text{val}})$, we compute the influence of each training example $(x_{i},y_{i})\in D_{\text{train}}$, where $D_{\text{train}}$ is the set of the training examples. The influence score $I((x_{i},y_{i}),(x_{\text{val}},y_{\text{val}}))$ quantifies the effect of $(x_{i},y_{i})$ on the loss function evaluated at $(x_{\text{val}},y_{\text{val}})$. Using these computed influence values, we select the top $k$ most influential demonstrations.
We compare two versions of computing the IF after extracting the gradient, DataInf Kwon et al. (2023) and TracIn Pruthi et al. (2020). DataInf is based on an easy-to-compute closed-form expression, leading to better computational and memory complexities than other IF computation methods. TracIn traces how the loss on the test point changes during the training process simply using an inner product of training set gradients. Since it does not compute the Hessian matrix, it is faster than DataInf, but at the cost of lower estimation performance.
Additionally, we compare the influence-only methods with well-performing approaches BertScore-Recall (BSR; Gupta et al. 2023a) and Cosine Similarity Reimers (2019). [^1] We use the implementation from Gupta et al. (2023a, b). These methods excel at capturing semantic similarity between validation and training examples. We also compare with a well performing sparse information retrieval algorithm, BM25 Jones et al. (2000).
Lastly, we combine the previously described approaches by implementing a two-stage selection process. First, we perform an initial pruning of the demonstration pool using either BSR or Cosine Similarity. Specifically, for a given number of desired demonstrations $k$, we prune the dataset to select $2k$ candidates from the original set of labeled examples $\{(x_{i},y_{i})\}_{i=1}^{N}$.
We then apply the IF-based methods to re-rank these remaining examples based on their influence on the validation loss. The final selection of $k$ in-context demonstrations is performed by selecting the top $k$ examples from the re-ranked subset.
3.2 Selecting Noisy Demonstrations
In this setting, we utilize IFs to identify noisy samples within the dataset. Formally, let $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ represent the training dataset. First, we employ IFs to prune the dataset by detecting and removing noisy examples, following which the top $k$ in-context demonstrations are selected using either BSR or Cosine Similarity Gupta et al. (2023a) [^2] We will refer to this approach as IF Pruning..
@@ 3.2 Selecting Noisy Demonstrations
where $I((x_{i},y_{i}),(x_{\text{val}},y_{\text{val}}))$ is the influence value and $S(x_{i},y_{i})$ represents either the BertScore Zhang et al. (2019) or Cosine Similarity for the training example $(x_{i},y_{i})$. The top $k$ examples with the highest combined scores are selected as demonstrations. [^3] We will refer to this approach as IF Averaging..
In this setting, we compute influence values using our surrogate model approach. In addition to using the IFs from DataInf Kwon et al. (2023), we also conduct influence experiments using the LiSSA method which is an iterative method to compute the inverse Hessian vector product Agarwal et al. (2017); Koh and Liang (2017).
@@ 4 Experiments
Here, we expand upon our experimental setup to conduct the experiments and analyze the results.
@@ 4.1 Experimental Setup
We discuss our, dataset details and model used to conduct the experiments.
@@
For Noisy Demonstrations, we employ the noisy dataset framework from Kwon et al. (2023). In their work, the four binary classification GLUE datasets Wang (2018) MRPC, QQP, QNLI, and SST2 are utilized. To simulate a scenario where a portion of the data samples are noisy, 20% of the training data samples are randomly selected and their labels are flipped. We use these noisy datasets as the candidate pool in our experiments and evaluate the ICL accuracy.
Base LLM. In Mixture of Tasks, for $k=3$ shots, we conduct ICL experiments on Llama-2-13b-chat Touvron et al. (2023), Mistral-7b-v0.3 Jiang et al. (2023) and Zephyr-7b-beta Tunstall et al. (2023). For $k=5$ shots we conduct experiments on Llama-2-13b-chat. We extend on the framework designed by Gupta et al. (2023a, b). The temperature is set to 0 for inference. For Noisy ICL, we conduct experiments on Llama-2-13b-chat. All of our experiments run on 8$\times$×NVIDIA RTX 6000 Ada GPUs.
@@ 4.2 Method and Baseline Configurations
Figure 2: Average performance of different demonstration selection methods across 3 LLMs for $k=3$ shots.
@@ 4.3 Results on Mixture of Tasks
We present the results on Mixture of Tasks in Figure 2, the full results for multiple number of shots and multiple LLMs can be seen in Appendix A.1. Additional results for the TracIn IF method are provided in Appendix A.2. Results for Pretrained Gradients combined with BertScore and Cosine Similarity are presented in Appendix A.3.
Combining Surrogate Model DataInf with BertScore results in the best performance. As we can clearly observe in Figure 2 and Table 8, Sur[D,BSR] method had the highest average performance across the tasks, in both 3 and 5 shots. This shows the benefit of combining IF with BertScore as performance increased by 0.56 in $k=3$ shots and by 1.52 in $k=5$ shots. The results also show that the maximal benefit of IF methods is gained in combination with the semantic similarity methods. This is due to the fact that IF can leverage the model’s inductive bias to re-rank the retrieved demonstrations effectively, but the initial $2k$ pruning via BSR is critical to shorten the candidate pool to demonstrations that are semantically relevant enough.
@@ 4.4 Results on Noisy ICL
In this section we provide analysis on the aforementioned General Noisy ICL setting. We also conduct two ablations on varying IF $\alpha$ weight in IF Averaging and varying the Noise level in the datasets.
@@
Method MRPC QNLI SST2 QQP PRU-0.2 Rand[Cos] 68.1 68.6 86.6 70.0 Sur[D,Cos] 67.7 67.2 86.2 70.0 Sur[L,Cos] 71.3 65.8 86.8 67.2 Rand[BSR] 69.1 72.2 85.4 69.8 Sur[D,BSR] 70.3 67.4 81.8 71.6 Sur[L,BSR] 70.1 66.4 85.2 70.8 PRU-0.3 Rand[Cos] 70.1 68.4 87.0 67.4 Sur[D,Cos] 68.8 69.2 84.6 70.6 Sur[L,Cos] 70.6 68.2 86.6 72.6 Rand[BSR] 70.8 65.8 86.0 70.8 Sur[D,BSR] 70.6 67.4 84.8 72.0 Sur[L,BSR] 69.4 68.2 87.4 68.4 AVG-0.4 Sur[D,Cos] 75.7 71.4 91.6 62.6 Sur[L,Cos] 73.3 67.2 90.6 75.2 Sur[D,BSR] 74.3 68.8 89.8 65.2 Sur[L,BSR] 56.6 72.8 78.0 73.0 AVG-0.6 Sur[D,Cos] 73.5 69.2 87.2 66.2 Sur[L,Cos] 72.1 68.6 84.6 70.8 Sur[D,BSR] 73.5 69.2 94.2 71.8 Sur[L,BSR] 72.3 65.8 94.2 75.4
Figure 4: ICL Accuracy for BertScore and Cosine Similarity with varying influence weight ($\alpha$) for MRPC 0.2
Varying IF $\alpha$ weight. We perform an analysis on the effect of varying alpha on the MRPC dataset with 20% noise added. We found that generally alpha ranges between 0.4 to 0.6 give the most optimal results as Figure 4 indicates. An explanation is that assigning higher weights to IF scores leads to missing out on semantically relevant samples and assigning higher weights to semantic selection methods leads to including several noisy samples.
Varying Noise level. We also test whether our method can perform well on varying noise levels in the dataset. To test this, we create 2 datasets of MRPC with 10% and 30% noise added. As seen in Table 3, in both the cases IF Averaging outperformed other baselines. With the DataInf configuration performing better for the 10% noise dataset and the LiSSA configuration performing better for the 30% noise dataset.

### v2

@@ Abstract
In this work, we introduce a novel paradigm for generalized In-Context Learning (ICL), termed Indirect In-Context Learning. In Indirect ICL, we explore demonstration selection strategies tailored for two distinct real-world scenarios: Mixture of Tasks and Noisy ICL. We systematically evaluate the effectiveness of Influence Functions (IFs) as a selection tool for these settings, highlighting the potential of IFs to better capture the informativeness of examples within the demonstration pool. For the Mixture of Tasks setting, demonstrations are drawn from 28 diverse tasks, including MMLU, BigBench, StrategyQA, and CommonsenseQA. We demonstrate that combining BertScore-Recall (BSR) with an IF surrogate model can further improve performance, leading to average absolute accuracy gains of 0.37% and 1.45% for 3-shot and 5-shot setups when compared to traditional ICL metrics. In the Noisy ICL setting, we examine scenarios where demonstrations might be mislabeled or have adversarial noise. Our experiments show that reweighting traditional ICL selectors (BSR and Cosine Similarity) with IF-based selectors boosts accuracy by an average of 2.90% for Cosine Similarity and 2.94% for BSR on noisy GLUE benchmarks. For the adversarial sub-setting, we show the utility of using IFs for task-agnostic demonstration selection for backdoor attack mitigation. Showing a 32.89% reduction in Attack Success Rate compared to task-aware methods. In sum, we propose a robust framework for demonstration selection that generalizes beyond traditional ICL, offering valuable insights into the role of IFs for Indirect ICL.
@@ 1 Introduction
Figure 1: Example showcasing demonstration selection for Indirect ICL using Influence Functions (IFs). Consider web corpora with many tasks (different from the end-task) and noisy data– Indirect ICL can be formalized as: Mixture of Tasks (Section 3.1) and Noisy (Section 3.2) ICL, respectively. In MoT, for a given target task (e.g. Medical Genetics), we first filter from this (indirect) pool of candidate demonstrations using BertScore and Cosine Similarity, then re-rank with IFs to select suitable demonstrations (e.g. High-School Biology). For Noisy ICL, we leverage IFs to filter out the Noisy ICL before conducting ICL with the remaining clean demonstrations.
Traditional ICL makes numerous assumptions that restrict its applicability to real-world problem domains. For instance, traditional ICL Min et al. (2021); Conneau (2019); Halder et al. (2020) assumes that demonstrations to be selected are directly and accurately annotated for the end-task. However, this is not always the case – for low-resource, sparse, or specialized domains, end-task information and labeled demonstrations might not be available. [^1] Consider the cases where we need to utilize ICL for diagnosing rare medical conditions, niche programming languages or indigenous spoken languages. Similarly, when LLMs are deployed as services, the user query or the end task itself could be unknown beforehand, let alone providing direct demonstrations at inference. [^2] Our proposed method can improve performance by selecting relevant demonstrations from a task agnostic pool of labeled data at test time. Thus, in this paper, we explore a more generalized setting for ICL, which we refer to as Indirect ICL.
In Indirect ICL, we aim to provide indirect (or incidental) supervision Yin et al. (2023); Li et al. (2024) by selecting demonstrations from a pool of examples where the majority are not directly suited to the end task due to severe distribution/covariate shifts. This includes selecting demonstrations from a pool that predominantly consists of demonstrations belonging to other tasks, with few demonstrations from the end task possibly included. Additionally, the demonstration set may be mislabeled by humans Yan et al. (2014); Zhu et al. (2022) or LLMs Wu et al. (2023). Since the effectiveness of ICL heavily relies on the quality of demonstrations selected Kossen et al. (2024); Wu et al. (2022); Wang et al. (2024), selecting the most helpful indirect demonstrations becomes imperative in these situations.
@@
Traditional (direct) ICL methods that use metrics such as BertScore-Recall (BSR; Gupta et al. 2023a) and cosine similarity Reimers (2019) inherently rely on the semantic similarity between demonstrations and test samples. In this paper, we posit that IFs can be a reasonable measure of affinity between the end task and any (indirect) demonstrations. We show that it is practical to use IFs to identify candidate demonstrations that represent a close inductive bias with the end-task, and utilize this information for highly accurate demonstration selection in the challenging Indirect ICL setting. As our experiments and results will demonstrate, this is indeed the case, and we find that IFs can aid in improved performance when simple semantic similarity is insufficient for demonstration selection. We provide additional examples of practical applications of Indirect ICL in Appendix A.
@@
  • We formalize a new and general paradigm for ICL, namely Indirect In-Context Learning, where we benchmark demonstration selection for two distinct and real-world settings: (a) Mixture of Tasks and (b) Noisy ICL. This novel paradigm with two settings is ubiquitous in the real world, and has yet been overlooked by existing research in ICL that assumes the availability of direct supervision.
  • We propose utilizing Influence Functions (IFs) as an effective approach for demonstration selection in generalized ICL settings, leveraging their capacity to exploit the task inductive bias of models to enhance selection quality. We also examine multiple influence functions for Indirect ICL and conduct an extensive analysis on their benefits in this setting.
  • For Mixture of Tasks, combining an IF Surrogate model with BertScore-Recall (BSR) can lead to a 0.37% and 1.45% average absolute increase in performance for $k=3$ and $k=5$ shots compared to the best performing traditional ICL metric.
  • For Noisy ICL, we observe that undertaking a weighted average selection using traditional ICL selectors (BSR and Cosine Similarity) and IF based selectors increases the absolute accuracy on average by 2.90% for Cosine and 2.94% for BSR for mislabeled samples. IFs can also lead to a 32.89% reduction in the Attack Success Rate for task-agnostic demonstration selection for backdoor attack mitigation.
@@ 2.1 Traditional In-Context Learning
2.2 Indirect In-Context Learning
@@
Noisy ICL. To further generalize the problem of Indirect ICL, we also consider noisy supervision that is likely existing in the pool of demonstrations. Formally, let $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ represent the training dataset, where $x_{i}\in X$ is the input and $y_{i}\in Y$ is the corresponding binary label. We randomly select a percentage of the data points from $D$ and flip their labels. Once the noisy dataset is generated, we use it for ICL. Formally, given the noisy set of input-output pairs $\{(x_{i},y_{i})\}_{i=1}^{k}$ and a test input $x_{\text{test}}$, the goal is to conditionally generate the test output $y_{\text{test}}$ based on the noisy training data.
@@ 3.1 Selecting within Mixture of Tasks
  • A pretrained-gradient based method where the samples are passed through the LLM itself. We then compute IFs using the extracted gradients.
Formally, for each validation example $(x_{\text{val}},y_{\text{val}})$, we compute the influence of each training example $(x_{i},y_{i})\in D_{\text{train}}$, where $D_{\text{train}}$ is the set of the training examples. The influence score $I((x_{i},y_{i}),(x_{\text{val}},y_{\text{val}}))$ quantifies the effect of $(x_{i},y_{i})$ on the loss function evaluated at $(x_{\text{val}},y_{\text{val}})$ [^3] We use a negative log-likelihood function as the loss function. Using these computed influence values, we select the top $k$ samples with the highest IF scores.
We compare two versions of computing the IF after extracting the gradient, DataInf Kwon et al. (2023) and TracIn Pruthi et al. (2020). DataInf uses an easy-to-compute closed-form expression, leading to better computational and memory complexities than other IF methods, more details in Appendix B. TracIn traces how the loss on the test point changes during the training process simply using an inner product of training and validation set gradients. Since it does not compute the Hessian matrix, it is faster than DataInf, but at the cost of lower estimation performance.
Additionally, we compare the influence-only methods described above with strong ICL baselines BertScore-Recall (BSR; Gupta et al. 2023a) and Cosine Similarity Reimers (2019). [^4] We use the implementation from Gupta et al. (2023a, b). These methods excel at capturing semantic similarity between validation and training examples. We also compare with a performant sparse information retrieval baseline algorithm, BM25 Jones et al. (2000).
Lastly, we combine the previously described approaches by implementing a two-stage selection process. First, we perform an initial pruning of the demonstration pool using either BSR or Cosine Similarity. Specifically, for a given number of desired demonstrations $k$, we prune the dataset to select $2k$ candidates from the original set of labeled examples $\{(x_{i},y_{i})\}_{i=1}^{N}$. We then apply the IF-based methods to re-rank these remaining examples based on their influence scores. The final selection of $k$ in-context demonstrations is performed by selecting the top $k$ examples from the re-ranked subset.
3.2 Selecting Noisy ICL
In this setting, we utilize IFs to identify noisy mislabeled samples within the dataset. Formally, let $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ represent the training dataset. First, we employ IFs to prune the dataset by detecting and removing noisy examples, following which the top $k$ in-context demonstrations are selected using either BSR or Cosine Similarity. We will refer to this approach as IF Pruning.
@@ 3.2 Selecting Noisy ICL
where $I((x_{i},y_{i}),(x_{\text{val}},y_{\text{val}}))$ is the influence value and $S(x_{i},y_{i})$ represents either the BertScore Zhang et al. (2019) or Cosine Similarity for the training example $(x_{i},y_{i})$. The top $k$ examples with the highest combined scores are selected as demonstrations. We will refer to this approach as IF Averaging.
In this setting, we compute influence values using our surrogate model approach. In addition to using DataInf, we also conduct influence experiments using the LiSSA IF method which is a second-order method to compute the inverse Hessian vector product Agarwal et al. (2017); Koh and Liang (2017). Although LiSSA is generally computationally expensive Kwon et al. (2023), we prioritize it over TracIn owing to its greater performance in detecting mislabeled samples, as the computational overhead is incurred only once in this setting [^5] We use different algorithms since Noisy ICL computes IF scores using all validation samples, enabling detection of noisy training samples. In contrast, MoT requires sample-specific IF scores, necessitating dynamic, per-instance computation. .
@@ 4 Experiments
Here, we expand upon our experimental setup to conduct the experiments and analyze the results. Overall, we find that combining DataInf with BSR for demonstration selection consistently improves accuracy across both the Mixture of Tasks and Noisy ICL settings within the Indirect ICL framework. Additionally, the surrogate model approach outperforms the pretrained gradient approach.
@@ 4.1 Experimental Setup
We discuss our dataset details and model used to conduct the experiments.
@@
For Noisy ICL, we employ the noisy dataset framework from Kwon et al. (2023). In their work, the four binary classification GLUE datasets Wang (2018) MRPC, QQP, QNLI, and SST2 are utilized. To simulate a scenario where a portion of the data samples are noisy, 20% of the training data samples are randomly selected and their labels are flipped. We use these noisy datasets as the candidate pool in our experiments and evaluate the ICL accuracy.
Base LLM. In Mixture of Tasks, for $k=3$ shots, we conduct ICL experiments on Llama-2-13b-chat Touvron et al. (2023), Mistral-7b-v0.3 Jiang et al. (2023), Zephyr-7b-beta Tunstall et al. (2023) and Qwen2.5-3b Team (2024). For $k=5$ shots we conduct experiments on Llama-2-13b-chat. We extend on the framework designed by Gupta et al. (2023a, b). The temperature is set to 0 for inference. For Noisy ICL, we conduct experiments on Llama-2-13b-chat for mislabeled data detection and Llama-3-8b for backdoor defense. All of our experiments run on 8$\times$×NVIDIA RTX 6000 Ada GPUs.
@@ 4.2 Method and Baseline Configurations
Figure 2: Average performance of different demonstration selection methods across Llama2-13b-chat, Zephyr-7b-beta and Mistral-7b-v0.3 for $k=3$ shots.
@@ 4.3 Results on Mixture of Tasks
We present the results on Mixture of Tasks in Figure 2. Additional results for varying the number of shots ($k$) and multiple LLMs are provided in Appendix C.1. Further, results for the alternative TracIn IF method are provided in Appendix C.2. Results for Pretrained Gradients combined with BertScore and Cosine Similarity are presented in Appendix C.3. We also present results on using DeBERTa He et al. (2021a, b) as an alternative surrogate model in Appendix C.4. Finally, we present results on Qwen2.5-3b Team (2024) in Appendix C.5.
Combining Surrogate Model DataInf with BertScore results in the best performance. As can be observed in Figure 2 and Table 6, the Sur[D,BSR] method has the highest average performance across the tasks, in both 3 and 5 shots. This shows the benefit of combining IF with BertScore as performance increased by 0.56 in $k=3$ shots and by 1.52 in $k=5$ shots. The results also show that the maximal benefit of IF methods is gained in combination with the semantic similarity methods. This is due to the fact that IF can leverage the model’s inductive bias to re-rank the retrieved demonstrations effectively, but the initial $2k$ pruning via BSR is critical to shorten the candidate pool to demonstrations that are semantically relevant enough. However, it is important to note that, given the inclusion of only three shots in the prompt, where the overwhelming majority of demonstrations are unrelated to the test task, achieving significant improvements remains challenging. We provide an ablation diving deeper into why combining IF with semantic similarity methods works well in Appendix C.6.
@@
Qualitative Analysis. Finally, to understand the unique benefits provided by IFs, we present a qualitative analysis examining the types of shots selected by our method in Appendix C.7. We see that even though BSR selects more semantically relevant samples, Sur[D,BSR] shots assist in guiding the model toward the correct answer by providing examples that promote more structured reasoning.
@@ 4.4 Results on Noisy ICL
In this section, we present results for the mislabeled data setting under Noisy ICL. Additionally, we report results for adversarial noise, specifically a backdoor defense strategy, in Appendix H.
@@
We further provide results on varying the noise levels in the datasets in D.1, varying the hyperparameters we tested in D.2 and an experiment analyzing the effectiveness of IF’s in detecting Noisy ICL in D.3.